# A deconstruction of methods to derive one-point lensing statistics

### v1

@@ Abstract
Gravitational lensing is a crucial tool for exploring cosmic phenomena, providing insights into galaxy clustering, dark matter, and dark energy. Given the substantial computational demands of $N$-body simulations, approximate methods like PINOCCHIO and turboGL offer viable alternatives for simulating lensing probability density functions (PDFs). This paper evaluates these methods in contexts where baryonic effects are negligible, focusing on dark matter-dominated models and assessing their effectiveness across both weak and strong lensing regimes. Our comparative analysis reveals that these methods are particularly effective for applications involving electromagnetic and gravitational wave point sources, where strong lensing events are infrequent. Both PINOCCHIO and turboGL perform well in modeling the weak-lensing region influenced by mildly nonlinear structures. However, they lose accuracy in capturing small-scale nonlinear matter fields, owing to oversimplified assumptions about internal halo structures and reliance on perturbation theory. The analysis shows that PINOCCHIO achieves an 8–15% agreement with $N$-body simulations for the second-to-fourth moments of lensing PDFs. These findings aim to inform future studies on gravitational lensing of point sources, which are increasingly relevant with upcoming supernova and gravitational wave datasets.
@@ 1 Introduction
Electromagnetic and gravitational wave point sources, such as supernovae and compact binary coalescences, have proven to be invaluable for studying the universe’s evolution across extensive redshift ranges. Upcoming observatories like the Euclid mission, [^2] https://www.euclid-ec.org/ the Nancy Grace Roman Space Telescope, [^3] https://roman.gsfc.nasa.gov/ and the Vera C. Rubin Observatory [^4] https://www.lsst.org/ are set to accurately observe tens of thousands of supernovae up to a redshift of $z\sim 2$. Simultaneously, third-generation gravitational wave detectors such as the Einstein Telescope [^5] https://www.et-gw.eu and Cosmic Explorer [^6] https://cosmicexplorer.org will detect hundreds of thousands of gravitational events up to $z\sim 5$ and beyond. To harness this wealth of data effectively, it is crucial to meticulously model all potential systematic effects, with gravitational lensing being a significant concern, particularly for sources at $z\gtrsim 1$, as it strongly influences distance measurements.
Gravitational lensing affects electromagnetic and gravitational waves as they traverse our inhomogeneous Universe. This phenomenon describes the deflections that waves undergo when interacting with the matter density field, causing distant sources to appear magnified or demagnified. Such deflections can lead to significant distance measurement errors, modifying the intrinsic distribution of astronomical sources and introducing variance and non-Gaussianities. Furthermore, this deflections also gives rise to distortions in the shapes of the background galaxies. A gravitational lens induces a correlation on galaxy image orientations which can be measured statically and reveal the properties and evolution of large-scale structures. This is known as cosmic shear (see e.g. Bartelmann and Schneider, 2001; Kilbinger, 2015, for a review).
@@
Accurate modeling of gravitational lensing effects on the Hubble diagram residuals can yield significant cosmological insights. By meticulously addressing the systematic effects of lensing, researchers can effectively transform what appears as noise into a valuable signal. Numerous studies (Quartin et al., 2014; Castro and Quartin, 2014; Castro et al., 2016b, a; Macaulay et al., 2017, 2020; Scovacricchi et al., 2017) have investigated the correlation between the Hubble diagram residuals and lensing signals, examining how these relationships depend on the background cosmology and the perturbations in intervening matter. These efforts have demonstrated the potential to achieve a precision of 0.6% for the amplitude of the power spectrum $\sigma_{8}$ and 7% for the growth rate index $\gamma$ in forthcoming surveys (Amendola et al., 2015).
The lensing probability distribution function (PDF) is the fundamental one-point statistic for describing the lensing of point sources. The most direct and precise method for determining the lensing PDF involves applying the ray-tracing technique in full hydrodynamical cosmological simulations (Castro et al., 2018). However, these simulations demand extensive computational resources. In regimes with negligible baryonic effects on lensing, $N$-body simulations using collisionless particles provide a sufficient alternative and significantly reduce computational demands. Despite the reliability of the results from $N$-body simulations (Takahashi et al., 2011; Hilbert et al., 2020), the computational requirements and storage capacities remain challenging.
@@
An even faster approach, which combines precision and speed, is to use approximate methods to describe the nonlinear evolution of perturbations (Manera et al., 2012; Kitaura and S., 2013; Tassev et al., 2013; Monaco, 2016). The PINpointing Orbit Crossing Collapsed HIerarchical Objects (PINOCCHIO) algorithm (Monaco et al., 2002; Taffoni et al., 2002; Munari et al., 2016) integrates Lagrangian perturbation theory (LPT) with a semi-analytical model to approximate the nonlinear hierarchical formation of structures. This method generates dark matter halo catalogs using a fraction of the resources needed for $N$-body simulations and achieves roughly 10% accuracy in clustering statistics (power spectrum, bispectrum, and two-point correlation function). For the weak-lensing regime, the stochastic gravitational lensing (sGL) approach provides an even faster alternative (Kainulainen and Marra, 2009, 2011a, 2011b). This method calculates the 1-halo term of the lensing PDF using stochastic configurations of inhomogeneities based on a halo model and subsequently incorporates the 2-halo term through convolution. A similar analytical method was recently proposed by Thiele et al. (2020).
@@
In this paper, we deconstruct the essential physics needed to calculate lensing PDFs accurately. We explore the effect of the 2-halo term, the influence of non-collapsed matter in the field, the accuracy of PINOCCHIO, and the degree to which baryonic effects can be overlooked. We specifically analyze four relevant lensing quantities in one-point statistics: the PDFs for convergence ($\kappa$), shear ($\gamma$), magnification ($\mu$), and magnitude magnification ($\Delta m$), the latter being particularly crucial for the statistics of lensing of standard candles like supernovae and gravitational waves. We will describe our methodology in Section 2, present our comparative results in Section 3, and summarize our findings in Section 4. In Appendix A, we revise the impact of the angular resolution on the lensing PDF.
2 Lensing Simulation Methods
@@
Hydro Ray-tracing on hydro-dynamical simulation PLC.
This study aims to assess the accuracy of various simulation methods in deriving the lensing Probability Density Function (PDF) on the regime in which baryonic physics is sub-dominant. The methods under investigation are summarized in Table 1, ranked by their complexity and the breadth of physical phenomena they encompass. For the PINOCCHIO and $N$-body simulations, we further evaluate the angular power spectrum and the generated lensing and density maps. The Hydro simulations are used simply to define the boundaries of reliability of the $N$-body results. For all methods, our fiducial model for the Universe is the $\Lambda$CDM model with the cosmological parameters set very close to the ones obtained by the combination of DES Y1 + Planck 2015 + JLA SNe + BAO (Abbott et al., 2018): $\Omega_{\rm m0}=0.301,\,\Omega_{\rm b0}=0.048,\,\Omega_{\Lambda 0}=0.699,\,h=0.682,\,\sigma_{8}=0.798$, and $n_{\rm s}=0.973$. We proceed to describe each method in detail.
@@ 2.1.1 The halo model
and $R$ is the halo radius. For the halo profile $\rho(r,M,z)$, we use the Navarro-Frenk-White (NFW) profile (Navarro et al., 1996). Modifications to the inner cuspy NFW profile by baryonic feedback are neglected.
@@ 2.1.2 Lensing convergence
Eq. (8) can be used to separate the contributions to the convergence due to small-scale inhomogeneities from the ones due to large-scale inhomogeneities so that, similarly to Eq. (1), it is:
@@
The third moment (skewness) can be obtained via the bispectrum (three-point correlation function), which should also contribute to the variance (although the contribution is expected to be small at these linear scales). Here, we take an alternative approach and model the full PDF relative to the 2-halo term as a log-normal distribution with zero mean and the variance given by Eq. (10), see Marra et al. (2013) for details. This is motivated by the well-known fact that the lensing PDF inherits the log-normal character of the density contrast PDF.
@@ 2.2 Light-cone construction and ray tracing
After creating the lens planes, the lensing quantities (convergence $\kappa$, shear $\gamma$, and magnification $\mu$) are computed using the ray-tracing technique, which integrates the non-linear matter on the unperturbed trajectory of the source plane to the observer. Castro et al. (2018) showed that the latter Born approximation yields accurate results for 1-point lensing statistics. In the thin lens approximation, i.e., disregarding the size of each lens, the convergence for a single lens plane is
@@
SMR uses FFTW to reconstruct the lensing potential from the convergence maps and fourth-order finite differences to calculate its shear components.
@@ 2.3 PINOCCHIO
Statistics Redshift Range
Convergence 1 [$\infty$- ∞, 0.38]
Convergence 2 $[-0.12,0.50]$
Convergence 3 $[-0.17,0.50]$
Convergence 5 $[-0.25,0.49]$
Shear 1 [$1.1\times 10^{-5}$, $5.8\times 10^{-2}$]
Shear 2 [$2.3\times 10^{-5}$, $1.3\times 10^{-1}$]
Shear 3 [$5.5\times 10^{-5}$, $1.5\times 10^{-1}$]
Shear 5 [$6.0\times 10^{-5}$, $1.9\times 10^{-1}$]
Magnification 1 $[0.93,2.3]$
Magnification 2 $[0.88,3.2]$
Magnification 3 $[0.83,2.4]$
Magnification 5 $[0.72,2.9]$
Magnitude 1 $[-0.90,0.08]$
Magnitude 2 $[-1.26,0.14]$
Magnitude 3 $[-0.95,0.20]$
Magnitude 5 $[-1.16,0.36]$
@@ 2.5 Hydrodynamical simulations
The effect of baryons on lensing probability density functions and the convergence power spectrum $P_{\kappa}(\ell)$ was thoroughly examined by Castro et al. (2018) using hydrodynamic simulations. This study specifically analyzed the multipole range of $P_{\kappa}(\ell)$ and the value ranges for $\kappa$, $\gamma$, and $\mu$ across various redshifts, identifying thresholds below which the baryonic contributions are minimal. Given the challenges associated with extracting cosmological information in scenarios where baryonic effects—such as AGN and stellar feedback, star formation, magnetic fields, and thermal conduction—are significant, our analysis adopts a simplified approach. We thus avoid the complexities of expensive hydrodynamical simulations by focusing on the dark matter-dominated regime, ensuring that baryonic effects remain below the 50% threshold. Although more stringent thresholds could be employed, it is pertinent to note that for many lensing statistics, like the moments of the PDFs (Marra et al., 2013), the discrepancies between DM-only and baryonic cases are minimal, as our DM model effectively captures the bulk of the distributions.
Table 3 displays the ranges within which the PDFs from the Hydro and DM versions of Box 3 (hr) of the Magneticum simulation suite [^11] http://www.magneticum.org differ by less than 50%, given a resolution of 1.76 arcsec. To avoid overcrowding in the plots, only the ranges corresponding to redshifts $z=1$ and $z=3$ will be presented.

### v2

@@ Abstract
Gravitational lensing is a crucial tool for exploring cosmic phenomena, providing insights into galaxy clustering, dark matter, and dark energy. Given the substantial computational demands of $N$-body simulations, approximate methods like PINOCCHIO and turboGL have been proposed as viable alternatives for simulating lensing probability density functions (PDFs). This paper evaluates these methods and their effectiveness across both weak and strong lensing regimes, with a focus in the context where baryonic effects are negligible. Our comparative analysis reveals that these methods are effective for applications where lensing is mild, such as the majority of sources of electromagnetic and gravitational waves. However, both PINOCCHIO and turboGL break down for large values of convergence and magnification due to their loss of accuracy in capturing small-scale nonlinear matter fields, owing to oversimplified assumptions about internal halo structures and reliance on perturbation theory. PINOCCHIO yields second-to-fourth moments of the lensing PDFs, which are 6–10% smaller than those resulting from $N$-body simulations in regimes where baryonic effects are minimal. These findings aim to inform future studies on gravitational lensing of point sources, which are increasingly relevant with upcoming supernova and gravitational wave datasets.
@@ 1 Introduction
Electromagnetic and gravitational wave (GW) point sources, such as supernovae (SN) and compact binary coalescences, have proven to be invaluable for studying the universe’s evolution across extensive redshift ranges. Upcoming observations such as those from the Euclid mission (Mellier et al., 2024), [^2] https://www.euclid-ec.org/ the Nancy Grace Roman Space Telescope, [^3] https://roman.gsfc.nasa.gov/ and the Vera C. Rubin Observatory [^4] https://www.lsst.org/ are set to accurately observe tens of thousands of supernovae up to a redshift of $z\sim 2$. Simultaneously, third-generation gravitational wave detectors such as the Einstein Telescope [^5] https://www.et-gw.eu and Cosmic Explorer [^6] https://cosmicexplorer.org will detect hundreds of thousands of gravitational events up to $z\sim 5$ and beyond. To harness this wealth of data effectively, it is crucial to meticulously model all potential systematic effects, with gravitational lensing being a significant concern, particularly for sources at $z\gtrsim 1$, as it strongly influences distance measurements.
Gravitational lensing affects electromagnetic and gravitational waves as they traverse our inhomogeneous Universe. This phenomenon describes the deflections that waves undergo when interacting with the matter density field, causing distant sources to appear magnified or demagnified. Such deflections can lead to significant distance measurement errors, modifying the intrinsic distribution of astronomical sources and introducing variance and non-Gaussianities. Furthermore, these deflections also give rise to distortions in the shapes of the background galaxies. A gravitational lens induces a correlation on galaxy image orientations, which can be measured statically and reveal the properties and evolution of large-scale structures. This is known as cosmic shear (see, e.g. Bartelmann and Schneider, 2001; Kilbinger, 2015, for a review).
@@
The lensing probability distribution function (PDF) is the fundamental one-point statistic for describing the lensing of point sources. After some pioneering work (e.g. Frieman 1996), various approaches have been followed so as to compute the lensing PDF relative to a given cosmology. A first approach (e.g. Munshi and Jain 2000; Valageas 2000; Wang et al. 2002; Das and Ostriker 2006) relates a “universal” form of the lensing PDF to the variance of the convergence, which is fixed by the amplitude $\sigma_{8}$ of the power spectrum. Moreover the coefficients of the proposed PDF may be trained on a grid of $N$-body simulations. Recently, using theoretical predictions from Large Deviation Theory to incorporate the non-Gaussian effects of the lensing PDF, Barthelemy et al. (2020) showed that it is possible to compute the lensing PDF consistently with results from simulations at $z=1$ and for an opening angle of 10 arcmin. Alternatively, one can build a model of the universe, e.g., using an $N$-body simulation, and directly compute the relative lensing PDF, usually through time-consuming ray-tracing techniques (e.g., Holz and Wald, 1998; Bergstrom et al., 2000; Holz and Linder, 2005; Hilbert et al., 2007, 2008; Takahashi et al., 2011a; Bolejko and Ferreira, 2012). The flexibility of this method is therefore penalized by the increased computational time.
@@
The most direct and precise method for determining the lensing PDF involves applying the ray-tracing technique in full hydrodynamical cosmological simulations (Castro et al., 2018). However, these simulations demand even more computational resources than their $N$-body counterpart. In regimes with negligible baryonic effects on lensing, $N$-body simulations using collisionless particles provide a sufficient alternative and significantly reduce computational demands. Despite the reliability of the results from $N$-body simulations (Takahashi et al., 2011b; Hilbert et al., 2020), the computational requirements and storage capacities remain challenging. An even faster approach, which searches for an optimum compromise between precision and speed, is to use approximate methods to describe the nonlinear evolution of perturbations (Manera et al., 2012; Kitaura and S., 2013; Tassev et al., 2013; Monaco, 2016). The PINpointing Orbit Crossing Collapsed HIerarchical Objects (PINOCCHIO) algorithm (Monaco et al., 2002; Taffoni et al., 2002; Munari et al., 2016) integrates Lagrangian perturbation theory (LPT) with a semi-analytical model to approximate the nonlinear hierarchical formation of structures. This method generates dark matter halo catalogs using a fraction of the resources needed for $N$-body simulations. It achieves roughly 10% accuracy in clustering statistics (power spectrum, bispectrum, and two-point correlation function) compared to the $N$-body solution. For the weak-lensing regime, the stochastic gravitational lensing (sGL) approach provides an even faster alternative (Kainulainen and Marra, 2009, 2011a, 2011b). This method calculates the 1-halo term of the lensing PDF using stochastic configurations of inhomogeneities based on a halo model and subsequently incorporates the 2-halo term through convolution. A similar analytical method, also based on combining the 1-halo with the 2-halo term, was recently proposed by Thiele et al. (2020).
The main importance of computing the lensing PDFs is understanding the extra scatter on future SN and GW measurements: an accurate modeling can yield significant cosmological insights. Since the final observed scatter on a standard candle Hubble diagram is a convolution of their intrinsic distribution with the lensing PDF at the observed redshift, one can transform what appears as noise in Hubble diagram residuals into a valuable signal. Numerous studies (Quartin et al., 2014; Castro and Quartin, 2014; Castro et al., 2016a, b; Macaulay et al., 2017; Scovacricchi et al., 2017; Macaulay et al., 2020; Shah et al., 2024) have investigated the correlation between these residuals and lensing signals, examining how these relationships depend on the background cosmology and the perturbations in the intervening matter. These efforts have demonstrated the potential to achieve a precision of 0.6% for the power spectrum amplitude $\sigma_{8}$ and 7% for the growth rate index $\gamma$ in forthcoming surveys (Amendola et al., 2015). One of the proposed methods to constrain cosmological parameters through the Hubble diagram residuals is the method of the moments (MeMo, see Quartin et al. 2014). [^7] See also Patton et al. (2017); Liu and Madhavacheril (2019); Ajani et al. (2023); Castiblanco et al. (2024). The MeMo relies on a comparison of the observed central moments of the residual magnitude distribution with the expectations from lensing simulations. Marra et al. (2013) computed fitting functions of the moments, but relied on the weak-lensing based turboGL and only for moderate redshifts. The next generations of SN surveys and GW detections will extend the observed redshift range, and to access all the cosmological information contained in these data more accurate tools that correctly describe the matter distribution in the universe are required. More recently, Boyle et al. (2021) showed that the lensing PDF provides, when combined with a Planck prior, more powerful constraints on the dark energy equation of state and neutrino mass than the two-point correlation function.
@@
In this paper, we deconstruct the essential physics needed to calculate lensing PDFs accurately. We explore the effect of the 2-halo term, the influence of non-collapsed matter in the field, the accuracy of PINOCCHIO, and the degree to which baryonic effects can be overlooked. We specifically analyze four relevant lensing quantities in one-point statistics: the PDFs for convergence ($\kappa$), shear ($\gamma$), magnification ($\mu$), and magnitude magnification ($\Delta m$), the latter being particularly crucial for the statistics of lensing of standard candles like SN and GW. We will describe our methodology in Section 2, present our comparative results in Section 3, and summarize our findings in Section 4. In Appendix A, we revise the impact of the angular resolution on the lensing PDF.
2 Methodology
@@
Hydro Ray-tracing on hydrodynamical simulation PLC.
This study aims to assess the accuracy of various simulation methods in deriving the lensing Probability Density Function (PDF) on the regime in which baryonic physics is sub-dominant. The methods under investigation are summarized in Table 1, ranked by their complexity and the breadth of physical phenomena they encompass. For the PINOCCHIO and $N$-body simulations, we further evaluate the angular power spectrum and the generated lensing and density maps. The Hydro simulations are used to define the boundaries of reliability of the $N$-body results. For all methods but the Hydro, our fiducial model for the Universe is the $\Lambda$CDM model with the cosmological parameters set very close to the ones obtained by the combination of DES Y1 + Planck 2015 + JLA SNe + BAO (Abbott et al., 2018): $\Omega_{\rm m0}=0.301,\,\Omega_{\rm b0}=0.048,\,\Omega_{\Lambda 0}=0.699,\,h=0.682,\,\sigma_{8}=0.798$, and $n_{\rm s}=0.973$. For the Hydro simulations, we rely on the lensing PDFs from Castro et al. (2018), based on the Magneticum simulations. [^8] http://www.magneticum.org/ Magneticum assumes the following background cosmology: $\Omega_{\rm m0}=0.272,\,\Omega_{\rm b0}=0.048,\,\Omega_{\Lambda 0}=0.728,\,h=0.704,\,\sigma_{8}=0.809$, and $n_{\rm s}=0.963$. We proceed to describe each method in detail in the order they appear in Table 1.
@@ 2.1.1 The halo model
and $R$ is the halo radius. For the halo profile $\rho(r,M,z)$, we use the Navarro-Frenk-White (NFW) profile (Navarro et al., 1996). Modifications to the inner cuspy NFW profile by baryonic feedback could be included if a suitable profile is provided (see, e.g., Di Cintio et al., 2014; Schneider and Teyssier, 2015).
@@ 2.1.2 Lensing convergence
Eq. (8) can be used to separate the contributions to the convergence due to small-scale inhomogeneities from the ones due to large-scale inhomogeneities so that, analogously to Eq. (1), it is: [^11] Note that the splitting is a priori different as power spectrum and lensing are affected by non-linearities in a different way.
@@
The third moment (skewness) can be obtained via the bispectrum (three-point correlation function), which should also contribute to the variance (although the contribution is expected to be small at these linear scales). Here, we take an alternative approach and model the full PDF relative to the 2-halo term as a shifted log-normal distribution with zero mean and the variance given by Eq. (10), see Marra et al. (2013) for details. This is motivated by the fact that the lensing PDF inherits – see Eq. (6) – the well-known log-normal character of the density contrast PDF (Coles and Jones, 1991; Kayo et al., 2001). As it can be seen by the example of Fig. 5, the final distribution is a minor modification to the corresponding Gaussian and captures the fact that, even in the linear regime, underdensities occupy more volume than overdensities, causing an intrinsic skewness in the distribution. See Barthelemy et al. (2024) for a model from first-principles that captures the skewness of the PDF in the quasi-linear regime.
@@ 2.2 Light-cone construction and ray tracing
After creating the lens planes, the lensing quantities (convergence, $\kappa$, shear, $\gamma$, and magnification, $\mu$) are computed using a ray-tracing technique that integrates the non-linear matter distribution along the unperturbed trajectory from the source plane to the observer. Castro et al. (2018) showed that, under the Born approximation, the 1-point lensing statistics deviates by less than 10% up to $z=5$ compared to the full solution within the 68% highest density interval. However, the tails of the distribution may be significantly affected. We independently validate the degree of validity of the Born approximation by comparing the results when using GLAMER deflection solver (Metcalf and Petkova, 2014; Petkova et al., 2014). Nevertheless, since our focus is on the relative differences between simulation methods, and we use the same ray-tracing algorithm for all simulations, the relative differences remain basically unaffected. In the thin lens approximation, i.e., disregarding the size of each lens, the convergence for a single lens plane is
@@
SMR uses FFT to reconstruct the lensing potential from the convergence maps and fourth-order finite differences to calculate its shear components. While tacitly assuming periodic boundary conditions when using FFT to reconstruct the shear maps is not formally correct, we have assessed its impact on the lensing PDF and found it to be completely subdominant compared to the differences between the simulation methods.
@@ 2.5 Hydrodynamical simulations
The effect of baryons on lensing probability density functions and the convergence power spectrum $P_{\kappa}(\ell)$ was thoroughly examined by Castro et al. (2018) using hydrodynamic simulations. This study specifically analyzed the multipole range of $P_{\kappa}(\ell)$ and the value ranges for $\kappa$, $\gamma$, and $\mu$ across various redshifts, identifying thresholds below which the baryonic contributions are minimal. Given the challenges associated with extracting cosmological information in scenarios where baryonic effects—such as AGN and stellar feedback, star formation, magnetic fields, and thermal conduction—are significant, our analysis adopts a simplified approach. We thus avoid the complexities of expensive hydrodynamical simulations by focusing on the dark matter-dominated regime, ensuring that baryonic effects remain below the 50% threshold. Table 3 displays the ranges within which the PDFs from the Hydro and DM versions of Box 3 (hr) of the Magneticum simulation suite differ by less than 50%, given a resolution of 1.76 arcsec. Although more stringent thresholds could be employed, as discussed in the next section, these values effectively limit our discussion to the bulk of the distribution, which is well described by a DM only model.
Statistics Redshift Range $\mathcal{L}_{1}$ [%]
Convergence 1 [$\infty$- ∞, 0.38] 1.3
Convergence 2 $[-0.12,0.50]$ 1.9
Convergence 3 $[-0.17,0.50]$ 2.5
Convergence 5 $[-0.25,0.49]$ 2.9
Shear 1 [$1.0\times 10^{-4}$, $3.5\times 10^{-2}$] 2.1
Shear 2 [$1.0\times 10^{-4}$, $6.0\times 10^{-2}$] 3.0
Shear 3 [$1.0\times 10^{-4}$, $8.0\times 10^{-2}$] 3.7
Shear 5 [$1.0\times 10^{-4}$, $1.0\times 10^{-1}$] 4.1
Magnification 1 $[0.93,2.3]$ 1.3
Magnification 2 $[0.88,2.3]$ 1.9
Magnification 3 $[0.83,2.4]$ 2.4
Magnification 5 $[0.72,2.9]$ 2.8
Magnitude 1 $[-1.17,0.11]$ 1.3
Magnitude 2 $[-1.34,0.24]$ 2.0
Magnitude 3 $[-1.52,0.32]$ 2.5
Magnitude 5 $[-1.52,0.46]$ 2.9
2.6 The $\mathcal{L}_{1}$ norm
The PDFs of Hydro and DM simulations differ more significantly on the tails; therefore, a local discrepancy of 50% usually corresponds to a much smaller global change of statistics. It is, hence, interesting to compute the weighted relative difference between a surrogate and reference PDFs:
$\mathcal{L}_{1}(a,b)\equiv\frac{1}{2}\int_{a}^{b}{\frac{\left|\frac{\textrm{d}P}{\textrm{d}x}_{\rm model}-\frac{\textrm{d}P}{\textrm{d}x}_{\rm surrogate}\right|}{\frac{\textrm{d}P}{\textrm{d}x}_{\rm model}}\,\frac{\textrm{d}P}{\textrm{d}x}_{\rm model}\textrm{d}x}\,,$ (29)
that tautologically reduces to
$\mathcal{L}_{1}(a,b)=\frac{1}{2}\int_{a}^{b}{\left|\frac{\textrm{d}P}{\textrm{d}x}_{\rm model}-\frac{\textrm{d}P}{\textrm{d}x}_{\rm surrogate}\right|\textrm{d}x}\,,$ (30)
which gives the cumulative absolute difference between the distributions. We refer to this statistic as the $\mathcal{L}_{1}$ norm, in analogy to the metric with the same name for distance between two vectors.
In particular, $\mathcal{L}_{1}(-\infty,\infty)$ is a simple scalar that quantifies the global concordance of two PDFs. We will refer to $\mathcal{L}_{1}(-\infty,\infty)$ as simply $\mathcal{L}_{1}$ in what follows. In the limit in which both PDFs have zero overlap, we get $\mathcal{L}_{1}=1=100\%$, which can be interpreted as 100% of the observed histogram bins of the surrogate PDF will be misplaced with respect to the model. Hence the prefactor of 1/2 in Eq. (30). The last column of Table 3 shows $\mathcal{L}_{1}(a,b)$, relative to the range given in the third column, between the surrogate (DM) and reference (Hydro) PDFs. We see that a local discrepancy of $50\%$ in the PDFs corresponds to a weighted relative difference of 2% for magnification.